OpenStudy (ohohaye):

Write the product in standard form (1+3i)(5-i)

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

the adult way is \[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\]

OpenStudy (misty1212):

or you can just multiply out then remember that \(i^2=-1\) when you get it

OpenStudy (anonymous):